# Egorov-Type Semiclassical Limits for Open Quantum Systems with a Bi-Lindblad Structure

### v1

@@ 5 An Euler-top-type Poisson–Lie pencil as a bi-Lindblad system
In this section we illustrate how the abstract framework developed above can be implemented in a concrete low-dimensional model of Euler-top type, inspired by the deformed Euler top of Marrero and collaborators [ballesteros2017poisson, G.I.M+2023]. Rather than reproducing every physical feature of the rigid-body Euler top, we focus on the structural ingredients that are essential for our construction: (i) a linear Poisson pencil of Poisson–Lie origin, (ii) a bi-Hamiltonian vector field admitting a Lenard–Magri hierarchy of commuting invariants, (iii) a homogeneous lift yielding a completely integrable contact system with nontrivial dissipated quantities, and (iv) an algebraic setting in which one can build Lindblad generators implementing pure dephasing relative to the quantum integrals. In this sense, the example below should be read as a deformed-Euler-top-type testbed tailored to the bi-Lindblad viewpoint.
@@
Such exact symplectic realizations of linear Poisson manifolds are classical and can be constructed explicitly; see, for instance, [C.F.M2021, Section 6.3]. In particular, for $\mathfrak{so}(3)^{*}\simeq\mathfrak{su}(2)^{*}$ one may take $S=N\times\mathbb{R}^{3}_{m}$, where $N$ is a $3$-manifold equipped with a coframe $\{\vartheta_{1},\vartheta_{2},\vartheta_{3}\}$ satisfying the Maurer–Cartan relations, and define
@@
and hence $F_{1}$ and $F_{2}$ are homogeneous of degree $2$. Let us define the functions $I_{a}^{\Sigma}\colon U\to\mathbb{R}$, defined on the open subset $U=S\setminus(N\times\{0\})$, given by
@@
Semiclassically, the joint spectral parameter $\nu=(\nu_{0},\nu_{1})$ labels, up to the usual Bohr–Sommerfeld quantization conditions (see e.g. [GuilleminSternberg1990]), the common level sets $\{I_{0}=\nu_{0},\ I_{1}=\nu_{1}\}$ of the underlying contact integrable system.
@@
while the diagonal blocks $\rho_{\nu\nu}$ are unaffected by the dissipator. Hence the off-diagonal coherences decay exponentially,
$\|\rho_{\nu\mu}(t)\|\;\lesssim\;\exp\!\Big(-\tfrac{1}{2}\,t\sum_{k=1}^{N}(\ell_{k,\nu}-\ell_{k,\mu})^{2}\Big)\,\|\rho_{\nu\mu}(0)\|,\qquad\nu\neq\mu,$
@@
In the present minimal Euler-top pencil, one may simply take $H_{0}^{\mathrm{eff}}=I_{0}$ and $H_{1}^{\mathrm{eff}}=I_{1}$, up to irrelevant affine reparametrizations. Let $\widehat{H}_{0,\hbar}$ and $\widehat{H}_{1,\hbar}$ be corresponding semiclassical quantizations.
@@
By construction, the Lindblad operators are of Weyl type and their principal symbols depend only on the homogeneous lifts $(I_{0}^{\Sigma},I_{1}^{\Sigma})$. Hence condition (3) of Theorem 4.24 is satisfied for both generators. Moreover, the commutation relations (5.8) imply that $\widehat{I}_{0,\hbar}$ and $\widehat{I}_{1,\hbar}$ are common quantum constants of motion for both dynamics. The semiclassical limit of the Hamiltonian commutator then yields the corresponding contact dynamics in the sense of (CC2).
Therefore, the families $\{\mathcal{L}^{(0)}_{\hbar}\}$ and $\{\mathcal{L}^{(1)}_{\hbar}\}$ form a bi-Lindblad pair in the sense of Definition 4.21.
@@
This provides a natural setting in which decoherence, integrability, and the semiclassical limit are aligned by construction rather than enforced by model-dependent assumptions: decoherence dynamically selects the very algebra that encodes the classical integrable geometry. $\diamond$
@@ 6 Conclusions
Conflict of Interest Statement.
@@
Data Availability Statement.

### v2

@@ 5 An Euler-top-type Poisson–Lie pencil as a bi-Lindblad system
In this section we illustrate how the abstract framework developed above can be implemented in a concrete low-dimensional model of Euler-top type, inspired by the deformed Euler top of Marrero and collaborators [3, 33]. Rather than reproducing every physical feature of the rigid-body Euler top, we focus on the structural ingredients that are essential for our construction: (i) a linear Poisson pencil of Poisson–Lie origin, (ii) a bi-Hamiltonian vector field admitting a Lenard–Magri hierarchy of commuting invariants, (iii) a homogeneous lift yielding a completely integrable contact system with nontrivial dissipated quantities, and (iv) an algebraic setting in which one can build Lindblad generators implementing pure dephasing relative to the quantum integrals. In this sense, the example below should be read as a deformed-Euler-top-type testbed tailored to the bi-Lindblad viewpoint.
@@
Such exact symplectic realizations of linear Poisson manifolds are classical and can be constructed explicitly; see, for instance, [12, Section 6.3]. In particular, for $\mathfrak{so}(3)^{*}\simeq\mathfrak{su}(2)^{*}$ one may take $S=N\times\mathbb{R}^{3}_{m}$, where $N$ is a $3$-manifold equipped with a coframe $\{\vartheta_{1},\vartheta_{2},\vartheta_{3}\}$ satisfying the Maurer–Cartan relations, and define
@@
and hence $F_{1}$ and $F_{2}$ are homogeneous of degree $2$, i.e., $\Delta F_{a}=2F_{a},\,a=0,1$. Let us define the functions $I_{a}^{\Sigma}\colon U\to\mathbb{R}$, defined on the open subset $U=S\setminus(N\times\{0\})$, given by
@@
Semiclassically, the joint spectral parameter $\nu=(\nu_{0},\nu_{1})$ labels, up to the usual Bohr–Sommerfeld quantization conditions (see e.g. [32]), the common level sets $\{I_{0}=\nu_{0},\ I_{1}=\nu_{1}\}$ of the underlying contact integrable system.
@@
while the diagonal blocks $\rho_{\nu\nu}$ are unaffected by the dissipator. Hence the off-diagonal coherences decay exponentially, that is, there exists a constant $C>0$ such that, for all $t\geq 0$ and $\nu\neq\mu$,
$\|\rho_{\nu\mu}(t)\|\;\leq\;C\,\exp\!\Big(-\tfrac{1}{2}\,t\sum_{k=1}^{N}(\ell_{k,\nu}-\ell_{k,\mu})^{2}\Big)\,\|\rho_{\nu\mu}(0)\|\,$,
@@
In the present minimal Euler-top pencil, one may simply take $H_{0}^{\mathrm{eff}}=I_{0}$ and $H_{1}^{\mathrm{eff}}=I_{1}$, up to irrelevant affine reparametrizations. Let $\widehat{H}_{0,\hbar}$ and $\widehat{H}_{1,\hbar}$ be the corresponding semiclassical quantizations.
@@
By construction, the Lindblad operators are of Weyl type and their principal symbols depend only on the homogeneous lifts $(I_{0}^{\Sigma},I_{1}^{\Sigma})$. Hence condition (3) of Theorem 4.24 is satisfied for both generators. Moreover, the commutation relations (5.8) imply that $\widehat{I}_{0,\hbar}$ and $\widehat{I}_{1,\hbar}$ are common quantum constants of motion for both dynamics. The semiclassical limit of the Hamiltonian commutator then yields the corresponding contact dynamics in the sense of (CC2). Therefore, the families $\{\mathcal{L}^{(0)}_{\hbar}\}$ and $\{\mathcal{L}^{(1)}_{\hbar}\}$ form a bi-Lindblad pair in the sense of Definition 4.21.
@@
This provides a natural setting in which decoherence, integrability, and the semiclassical limit are aligned by construction rather than enforced by model-dependent assumptions: decoherence dynamically selects the algebra that encodes the classical integrable geometry. $\diamond$
@@ 6 Conclusions
Declarations
Funding
L. Colombo acknowledges financial support from Grant PID2022-137909-NB-C22 funded by the Spanish Ministry of Science and Innovation.
Competing Interests
@@
Data Availability Statement